OpenStudy (jhhhhhiiiii):

Express the complex number in trigonometric form. -3i

OpenStudy (jhhhhhiiiii):

@baru can you help with this one and just one more?

OpenStudy (baru):

think of it as: 0+3i can you find its magnitude?

OpenStudy (jhhhhhiiiii):

3?

OpenStudy (baru):

yep, so that complex number in trignometric form looks this way: \[3(\cos \theta + i \sin \theta)\] so i'm just going to factor out 3 from 0 + 3i to get 3(0 + i) so now the problem reduces to finding \(\theta\) such that \(cos \theta=0\) and \(sin \theta = 1\) so what is the value of \(\theta\) ?

OpenStudy (jhhhhhiiiii):

I'm not sure..

OpenStudy (baru):

look it up here: http://www.mathwords.com/t/t_assets/t80.gif

OpenStudy (baru):

you just need to find \(\theta\) such that cos\(\theta\)=0

OpenStudy (jhhhhhiiiii):

0?

OpenStudy (baru):

is cos(0) = 0 ?

OpenStudy (jhhhhhiiiii):

yes

OpenStudy (baru):

http://www.mathwords.com/t/t_assets/t80.gif look it up in that link, what is the value of cos(0) ?

OpenStudy (jhhhhhiiiii):

It's 1?

OpenStudy (jhhhhhiiiii):

Idk I'm sorry I'm not good at this

OpenStudy (phi):

one way to find the angle is to plot the point (x is the real part and y is the complex part) in other words, think of 0-3i as (0, -3)